OpenStudy (anonymous):

How do I solve this equation: ln(3x-5)=ln11+ln2? Why is the answer x=9?

OpenStudy (anonymous):

start with \[\ln(3x-5)=\ln(22)\] then it should be clear

OpenStudy (anonymous):

do you know how I can solve, lnx=ln3-ln(x+5)?